OpenStudy (anonymous):

A stone is thrown vertically upward with a speed of 11.0 m/s from the edge of a cliff 66.0 m high. (a) How much later does it reach the bottom of the cliff? s (b) What is its speed just before hitting? m/s (c) What total distance did it travel? m

OpenStudy (anonymous):

use equations of constant acceleration displacement s = -66, u = 11, a = -9.8, t = ? s = ut + 0.5at^2 -66 = 11t - 4.9t^2 4.9t^2 - 11t - 66 = 0 t = 4.96 secs

OpenStudy (anonymous):

find time to reach greatest height v = 0, u = 11, a = -9.8, t =? o = 11 + -9.8 * t t = -11/-9.8 = 1.12 secs find v when stone hits ground: u = 0, t = 4.96-1.12 = 3.84, v = ?, a=9.8 v = 0 +9.8 * 3.84 = 37.63 m/s

OpenStudy (anonymous):

distance travelled ascending u=11, a=9.8,s =?,v = 0 v^2 = u^2 -2*9.8 * s 0 = 121 - 19.6 s s = 121/19.6 = 6.173m distance travelled when passing intitial postion = 2 * 6.173 = 12.35m total distance = 12.35 + 66 = 78.35 m

OpenStudy (anonymous):

thank you thank you thank you thank you. could you help me with my other problem too?

OpenStudy (anonymous):

i'll try